OpenStudy (anonymous):

Please help find the midpoint of the line segment with end points (-1, -2) and (5, -1)?

OpenStudy (amistre64):

midpoint is the average between the points; add them up and divide by 2

OpenStudy (anonymous):

Thanks!

OpenStudy (amistre64):

(-1, -2) (5, -1) -------- (4,-3) / 2 = (?,?)

OpenStudy (amistre64):

(4/2, -3/2) = (2, -3/2) is what i get

OpenStudy (anonymous):

Y

OpenStudy (anonymous):

Yes, I believe that's what I figured. Thanks :)

OpenStudy (amistre64):

lol, good :)